# Maintaining Plasticity in Deep Continual Learning

### v1

@@ 3 Robust Loss of Plasticity in Permuted MNIST
4 Loss of Plasticity With Different Activations in the Slowly Changing Regression Problem
@@
There remains the issue of the network’s activation function. We have been using ReLU, the most popular choice at present, but there are a number of other possibilities. For these experiments, we switched to an even smaller, more idealized problem. Slowly-Changing Regression is a computationally inexpensive problem where we can run a single experiment on a CPU core in 15 minutes, allowing us to do extensive studies. As the name suggests, this is a regression problem—meaning the labels are real numbers, with a square loss, rather than nominal values with a cross-entropy loss—and the nonstationarity is slow and continual rather than abrupt as in a switch from one task to another. In Slowly-Changing Regression, we study loss of plasticity for networks with six different activation functions: sigmoid, tanh, ELU (Clevert et al., 2016), leaky-ReLU (Maas et al., 2013), ReLU (Nair & Hinton, 2010), and Swish (Ramachandran et al., 2018).
@@
In Slowly-Changing Regression, the learner receives a sequence of examples. The input for each example is a binary vector of size $m+1$. The first $f$ bits in the input vector change slowly. After every $T$ example, one of the first $f$ bits is chosen randomly, and its value is flipped. These first $k$ bits remain fixed for the next $T$ examples. The parameter $T$ allows us to control the speed at which the input distribution changes. The next $m-f$ bits are randomly sampled for each example. Lastly, the $(m+1)$-th bit is a bias term with a constant value 1.
The target output is generated by running input through a neural network, which is set at the start of the experiment and kept fixed. As this network generates the target output and represents the desired solution, we call it the target network. The weights of the target networks are randomly chosen to be +1 or -1. The target network has one hidden layer with a Linear Threshold Unit (McCulloch & Pitts, 1943), or LTU, activation. The value of the $i$-th LTU is one if the input is above a threshold $\theta_{i}$, and zero otherwise. $\theta_{i}$ is set equal to $(m+1)\cdot\beta-S_{i}$, where $\beta\in[0,1]$ and $S_{i}$ is the number of input weights with negative value (Whitehead and Sutton, 1994). The details of the input and target function in the slowly-changing regression problem are also described in Figure 4.
The details of the specific instance of the slowly-changing regression problem we use in this paper and the learning network used to predict its output are listed in Table 4. Note that the target network is more complex than the learning network as the target network is wider, with 100 hidden units, while the learner has just five hidden units. Thus, because the input distribution changes every $T$ example and the target function is more complex than what the learner can represent, there is a need to track the best approximation.
@@
We applied learning networks with different activation functions to the slowly-changing regression. The learner used the backpropagation algorithm to learn the weights of the network. We used uniform Kaiming distribution (He et al., 2015) to initialize the learning network’s weights. The distribution is $U(-b,b)$ with bound, $b=gain*\sqrt{\frac{3}{num\_inputs}}$, where $gain$ is chosen such that the magnitude of inputs does not change across layers. For tanh, Sigmoid, ReLU, and Leaky-ReLU, $gain$ is 5/3, 1, $\sqrt{2}$, and $\sqrt{2/(1+\alpha^{2})}$ respectively. For ELU and Swish, we used $gain=\sqrt{2}$, as was done in the original works by Clevert et al. (2016) and Ramachandran et al. (2018).
@@
The results of the experiments are shown in Figure 5. We measured the squared error, the square of the difference between the true target and the prediction made by the learning network. In Figure 5, the squared error is presented in bins of 40k examples. This means that the first data point is the average squared error on the first 40k examples, the next is the average squared error on the next 40k examples, and so on. The shaded region in the Figure shows the standard error of the binned error.
Figure 5: Loss of plasticity is robust across different activations.
Figure 5 shows that in Slowly-changing regression, after performing well initially, the error increases for all step sizes and activations. For some activations like ReLU and tanh, loss of plasticity was severe, and the error increased to the level of the linear baseline. While for other activations like ELU, loss of plasticity was less severe, but still there was a significant increase in the error. These results mean that the loss of plasticity is not resolved by using different activations.
So far in the paper, we performed direct tests of loss of plasticity with backpropagation. Our experiments in Continual Imagenet, online Permuted MNIST, and Slowly changing regression show that loss of plasticity is a general phenomenon, and it can be catastrophic in some cases. It happens for a wide range of hyper-parameters, activation functions, rates of distribution change, and for both under and over-parameterized networks. Although continual learning has been studied for a long time from (Caruana, 1997; Ring, 1998; Parisi et al. (2019)), a direct test of the ability to maintain plasticity was missing. The experiments so far in the paper fill this gap and show that conventional backpropagation does not work in continual learning.
There are two major goals in continual learning: maintaining stability and maintaining plasticity. Maintaining stability is concerned with memorizing useful information, and maintaining plasticity is about finding new useful information when the data distribution changes. It is difficult for current deep-learning methods to remember useful information as they tend to forget previously learned information (McCloskey & Cohen, 1989; French, 1999; Parisi et al., 2019). Various ingenious methods have been proposed to tackle the forgetting issue (Kirkpatrick et al., 2017; Aljundi et al., 2019; Riemer et al., 2019; Javed & White, 2019; Veniat et al., 2021), but it remains an unsolved problem. We focused on continually finding useful information, not on remembering useful information. Our work on loss of plasticity is different but complementary to the work on maintaining stability.

### v3

@@ Appendix A Methods
Appendix B Loss of Plasticity With Different Activations in the Slowly Changing Regression Problem
@@
There remains the issue of the network’s activation function. We have been using ReLU, the most popular choice at present, but there are a number of other possibilities. For these experiments, we switched to an even smaller, more idealized problem. Slowly-Changing Regression is a computationally inexpensive problem where we can run a single experiment on a CPU core in 15 minutes, allowing us to do extensive studies. As its name suggests, this problem is a regression problem—meaning the labels are real numbers, with a squared loss, rather than nominal values with a cross-entropy loss—and the nonstationarity is slow and continual rather than abrupt as in a switch from one task to another. In Slowly-Changing Regression, we study loss of plasticity for networks with six popular activation functions: sigmoid, tanh, ELU clevert2016 , leaky-ReLU mass2013 , ReLU nair2010 , and Swish ramachandran2018 .
@@
In Slowly-Changing Regression, the learner receives a sequence of examples. The input for each example is a binary vector of size $m+1$. The input has $f$ slow-changing bits, $m-f$ random bits, and then one constant bit. The first $f$ bits in the input vector change slowly. After every $T$ examples, one of the first $f$ bits is chosen uniformly at random, and its value is flipped. These first $f$ bits remain fixed for the next $T$ examples. The parameter $T$ allows us to control the rate at which the input distribution changes. The next $m-f$ bits are randomly sampled for each example. Lastly, the $(m+1)$-th bit is a bias term with a constant value 1.
The target output is generated by running input through a neural network, which is set at the start of the experiment and kept fixed. As this network generates the target output and represents the desired solution, we call it the target network. The weights of the target networks are randomly chosen to be +1 or -1. The target network has one hidden layer with a Linear Threshold Unit mcculloch1943 , or LTU, activation. The value of the $i$-th LTU is one if the input is above a threshold $\theta_{i}$, and zero otherwise. $\theta_{i}$ is set equal to $(m+1)\cdot\beta-S_{i}$, where $\beta\in[0,1]$ and $S_{i}$ is the number of input weights with negative value sutton1993 . The details of the input and target function in the slowly-changing regression problem are also described in Figure 9.
The details of the specific instance of the slowly-changing regression problem we use in this paper and the learning network used to predict its output are listed in Table B. Note that the target network is more complex than the learning network as the target network is wider, with 100 hidden units, while the learner has just five hidden units. Thus, because the input distribution changes every $T$ example and the target function is more complex than what the learner can represent, there is a need to track the best approximation.
@@
Figure 10: Loss of plasticity is robust across different activations.
We applied learning networks with different activation functions to the slowly-changing regression. The learner used the backpropagation algorithm to learn the weights of the network. We used uniform Kaiming distribution he2015delving to initialize the learning network’s weights. The distribution is $U(-b,b)$ with bound, $b=gain*\sqrt{\frac{3}{num\_inputs}}$, where $gain$ is chosen such that the magnitude of inputs does not change across layers. For tanh, Sigmoid, ReLU, and Leaky-ReLU, $gain$ is 5/3, 1, $\sqrt{2}$, and $\sqrt{2/(1+\alpha^{2})}$ respectively. For ELU and Swish, we used $gain=\sqrt{2}$, as was done in the original papers clevert2016 ; ramachandran2018 .
@@
The results of the experiments are shown in Figure 10. We measured the squared error, the square of the difference between the true target and the prediction made by the learning network. In Figure 10c, the squared error is presented in bins of 40k examples. This means that the first data point is the average squared error on the first 40k examples, the next is the average squared error on the next 40k examples, and so on. The shaded region in the Figure shows the standard error of the binned error.
Figure 10 shows that in Slowly-changing regression, after performing well initially, the error increases for all step sizes and activations. For some activations like ReLU and tanh, loss of plasticity is severe, and the error increases to the level of the linear baseline. While for other activations like ELU, loss of plasticity is less severe, but still there is a significant loss of plasticity. These results mean that the loss of plasticity is not resolved by using commonly used activations.